OpenStudy (anonymous):

Find the general solution to the follow differential equation: f′′ − 6f′ + 9f = 20 − 12x + 9x^2.

OpenStudy (anonymous):

i understand how to get the e part i just need help with the end part!

OpenStudy (amistre64):

whats you characteristic equation to solve for your fs

OpenStudy (amistre64):

oh, the end can be tabled

OpenStudy (amistre64):

yp = ax^2 +bx +c and take derivatives and plug them into your diffyQ

OpenStudy (amistre64):

well, fp for this since they wanna be weird and all :)

OpenStudy (anonymous):

ahah wait what!

OpenStudy (anonymous):

what derivitives do i plug into where?

OpenStudy (amistre64):

the derivatives of the generic quadratic that matches up to the specific quadratic

OpenStudy (amistre64):

fp = ax^2 + bx + c is a generic match to 9x^2 -12x +20

OpenStudy (amistre64):

find fp' and fp'' and plug then into the diffQ to compare parts with

OpenStudy (anonymous):

so fp' = 18x - 12, fp'' = 18?

OpenStudy (amistre64):

fp′′ − 6fp′ + 9fp = 9x^2-12x+20 fp = ax^2 + bx + c fp' = 2ax + b fp'' = 2a

OpenStudy (amistre64):

fp′′ − 6fp′ + 9fp = 9x^2-12x+20 \[2a -6(2ax + b) +9(ax^2 + bx + c) = 9x^2-12x+20\]

OpenStudy (anonymous):

ohhh ok.

OpenStudy (anonymous):

then do i factor an x out of like termS? or?

OpenStudy (amistre64):

expand it all; gather up your like terms; and compare them to their respective parts from the right side

OpenStudy (amistre64):

9ax^2 = 9x^2; a=1 that one an easy one

OpenStudy (anonymous):

so a = 1, b = -4/3, c=20/9 ?

OpenStudy (amistre64):

9bx + 12ax = -12x (9b +12)x = -12x 9b = -24; b=-24/9 = -12/6 = -4/3 yeah, stuff like that

OpenStudy (amistre64):

fp = ax^2 +bx + c we found a,b and c fp = x^2 -4x/3 +20/9

OpenStudy (amistre64):

you say you found the homogenous part; so add them up f = fh + fp

OpenStudy (anonymous):

ok so f(x) = Ae^3x + Bxe^3x + x^2 -4x/3 +20/9

OpenStudy (amistre64):

yep, as long as your fh is good ;)

OpenStudy (amistre64):

there might be an issue with your x terms tho; i recall someting about something that i forgot a while ago that might make that one an issue

OpenStudy (anonymous):

uh oh, whys that!

OpenStudy (amistre64):

something to do with keeping all the parts linearly independant

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%27%27-6y%27%2B9y%3D9x%5E2-12x%2B20 yeah, looks like that might be an issue since the solutions dont match up for us

OpenStudy (anonymous):

oh no so what did i do wrong ahha

OpenStudy (amistre64):

try it again but we have to bump up the xparts; x^2(ax^2+bx+c) ax^4+bx^3+cx^2 would seem to me to cover it; but thats just me trying to remember this

OpenStudy (anonymous):

hmmm ok

OpenStudy (anonymous):

i dont think i'm getting it still ..

OpenStudy (amistre64):

i checked the wolfs steps and they did what we did; i think i typed it in right :) double check the maths from the start of it

OpenStudy (amistre64):

2a -12ax -6b 9ax^2+9bx+9c -------------- 9x^2−12x+20 a=1 2 -12x -6b 9x^2+9bx+9c -------------- 9x^2−12x+20 -12+9b=-12 ; found it ... b=0 2 -12x - 9x^2+ +9c -------------- 9x^2−12x+20 2+9c = 20 ; c=2

OpenStudy (amistre64):

fp = x^2 + 2 :D